# Large Language Model Supply Chain: A Research Agenda

### v2

@@ Abstract.
The rapid advancement of Large Language Models (LLMs) has revolutionized artificial intelligence, introducing unprecedented capabilities in natural language processing and multimodal content generation. However, the increasing complexity and scale of these models have given rise to a multifaceted supply chain that presents unique challenges across infrastructure, foundation models, and downstream applications. This paper provides a comprehensive research agenda of the LLM supply chain, offering a structured approach to identify critical challenges and opportunities through the dual lenses of Software Engineering (SE) and Security & Privacy (S&P). We begin by establishing a clear definition of the LLM supply chain, encompassing its components and dependencies. We then analyze each layer of the supply chain, presenting a vision for robust and secure LLM development, reviewing the current state of practices and technologies, and identifying key challenges and research opportunities. This work aims to bridge the existing research gap in systematically understanding the multifaceted issues within the LLM supply chain, offering valuable insights to guide future efforts in this rapidly evolving domain.
@@ 1. Introduction
The landscape of artificial intelligence (AI) has been dramatically transformed by the rapid advancements in pre-trained Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs)  [^1] For simplicity in this text, both pre-trained LLMs and MLLMs will be collectively referred to as LLMs, and their supply chains will be referred to as the LLM supply chain in the subsequent sections.. Models such as GPT-4 (Achiam et al., 2023), Gemini (Team et al., 2023), and LLaMA (Touvron et al., 2023) have pushed the boundaries of what machines can understand and generate, exhibiting remarkable capabilities in processing and producing human-like text and multimodal content. These sophisticated models are trained on vast and diverse datasets, enabling them to perform a wide array of tasks ranging from natural language understanding and translation to image captioning and content creation.
As the adoption of LLMs continues to expand, the necessity for a robust and efficient supply chain to support their development, deployment, and maintenance becomes increasingly apparent. The LLM supply chain encompasses the lifecycle of these models, involving multiple stages such as data collection, model training, optimization, deployment, and continuous updating. This complex ecosystem requires seamless collaboration among various stakeholders, including model developers who design and refine architectures, data providers who supply necessary training data, infrastructure engineers who manage computational resources, and end-users who interact with final applications. Each stakeholder plays a critical role in ensuring the functionality and reliability of LLMs, and the interdependencies among them highlight the need for coordinated efforts to address challenges inherent in the supply chain.
@@
Current State & Challenges. Despite the significant progress and transformative potential of LLMs, the rapidly evolving landscape of their development and deployment presents many challenges across the entire supply chain. These challenges span multiple fileds, including security, privacy, software engineering, systems, networking, data science, and user experience design. In this paper, we focus specifically on Software Engineering (SE) and Security & Privacy (S&P). From the SE perspective, the unique characteristics of LLMs introduce a host of challenges that extend beyond traditional software development practices. The sheer scale and complexity of these models necessitate novel approaches to system design, development, and maintenance. Key challenges include model versioning and reproducibility (Ajibode et al., 2024), dependency management (Jiang et al., 2023), adapting continuous integration and deployment (CI/CD) practices for LLMs (Chen, 2024), comprehensive testing and quality assurance (Chang et al., 2024; Sun et al., 2024c), model interpretability and explainability (Singh et al., 2024), technical debt management (Bill Doerrfeld, 2024), and navigating the complex landscape of open-source licenses (Charlie Hull, 2023; Sriram Parthasarathy, 2024). From the S&P perspective, vulnerabilities have been identified at various layers of the LLM supply chain. These range from flaws in LLM development toolchains at the infrastructure layer (iumiro, 2023; Tom Bonner, 2023; Patrick Peng, 2024; Smartkeyss, 2024) to vulnerabilities (Eoin Wickens, Kasimir Schulz, 2024) or code poisoning (Zhao et al., 2024c) attacks in model hubs, and extend to remote code execution risks in downstream applications (Liu et al., 2023b). The pervasive nature of these vulnerabilities highlights the need for comprehensive security measures throughout the LLM supply chain. The potential for cascading effects due to model reuse and the difficulty in isolating and mitigating issues in complex LLM-based systems further exacerbate these challenges, often outstripping the establishment of robust security measures and software engineering practices.
@@
Research Gap & Our Work. Addressing these challenges necessitates a multidisciplinary approach that draws insights from fields such as software engineering, system architecture, security, and data governance. However, there exists a significant research gap in systematically tackling these multifaceted issues within the context of the LLM supply chain. To bridge this gap, this paper presents a comprehensive overview of the LLM supply chain, proposing a structured approach to analyze and address its challenges across three key layers: infrastructure, foundation models, and downstream applications. We begin by establishing a clear definition of the LLM supply chain, encompassing its components and dependencies. Building upon this foundation, we examine each layer through the dual lenses of SE and S&P. For each layer, we present a vision for the robust and secure LLM supply chain, review the current state of practices and technologies, and identify key challenges and research opportunities. In the infrastructure layer, we explore issues related to computational resources, development tools, and datasets. For foundation models, we delve into the complexities of model training, testing, sharing, and maintenance. The downstream applications layer focuses on the integration and deployment of LLM-based systems in real-world scenarios. Through this comprehensive examination, we propose a research agenda that not only identifies current challenges but also anticipates future developments in the field. Our goal is to offer insights that can guide future efforts in this rapidly evolving domain, fostering innovation while upholding ethical standards and ensuring the responsible advancement of LLM technologies.
@@
  • Comprehensive Overview of the LLM Supply Chain: The paper presents an in-depth examination of the entire LLM supply chain, highlighting its three core layers: the model infrastructure layer (including computational resources, datasets, and toolchains for training, optimization, and deployment), the foundation model layer (encompassing training, testing, releasing, and ongoing maintenance), and the downstream application ecosystem (facilitating the integration of pre-trained models into various intelligent applications).
@@
  • Proposal of A Research Agenda: Based on the systematic review, this paper proposes a forward-looking research agenda that addresses current challenges and anticipates future developments in the LLM supply chain. This agenda is designed to guide future efforts in the rapidly evolving domain of the LLM supply chain by suggesting multidisciplinary approaches that draw from SE and S&P. The goal is to foster advancements that maximize the transformative potential of LLMs while adhering to ethical standards and promoting sustainability.
@@ 2. Definitions and Significance
In this section, we start by defining the LLM supply chain (§ 2.1). We then explore three key aspects that highlight its significance: the introduction of novel challenges by LLMs beyond traditional open-source software (§ 2.2), the increased complexity of managing LLMs compared to conventional machine learning (ML) and deep learning (DL) models (§ 2.3), and the broader societal implications of LLMs, emphasizing the need for responsible AI practices within their supply chains (§ 2.4). These elements collectively underscore the urgent need for a comprehensive research agenda focused on the LLM supply chain.
@@ 2.1. Definitions
Building upon the concept of Open Source Software (OSS) supply chains (Wu and Lin, 2001; Scacchi et al., 2006; Ladisa et al., 2022), we define the LLM supply chain as a complex network system $S=(C,D)$, where $C$ represents the set of components and $D$ represents the set of dependencies. This system involves all upstream data providers, model development communities, model repositories, distribution platforms, and appmarkets, as well as data scientists, researchers, engineers, maintainers, and end-users in the process of developing, distributing, and deploying pre-trained models. As illustrated in Figure 1, the LLM supply chain comprises five key components $C=\{R,T,D,M,A\}$: computational resources ($R$), tools ($T$), datasets ($D$), models ($M$), and applications ($A$). The first three components ($R$, $T$, and $D$) collectively form the infrastructure of the LLM supply chain.
@@
  • LLM Toolchain ($T$): Encompasses development toolchains, machine learning frameworks, version control systems, and continuous integration tools.
@@
  • Datasets ($D$): Comprises large-scale text corpora, specialized domain datasets, and multi-modal data collections.
@@
  • Models ($M$): Represents the various stages of LLM development, including architecture design, pre-training, fine-tuning, testing, evaluation, deployment, and maintenance.
@@
  • Applications ($A$): Includes downstream LLM-based applications such as chatbots, intelligent agents, and domain-specific LLM solutions.
Figure 1. Definition and Each Component of the LLM Supply Chain.
Within this supply chain, there exist multiple complex dependencies $D={d_{ij}|i,j\in C}$, including:
  • $d_{RR}$: Interdependencies between different computational resources, potentially affecting performance and security.
  • $d_{RT}$: Compatibility and optimization between computational resources and development tools.
  • $d_{RD}$: Requirements of computational resources for efficient dataset processing and storage.
  • $d_{RM}$: Computational resource requirements for model training, fine-tuning, and inference.
@@
  • $d_{TT}$: Interoperability and version compatibility between different development tools.
  • $d_{TD}$: Effects of data processing tools on dataset quality and integrity.
  • $d_{TM}$: Potential introduction of vulnerabilities or optimizations through the development process.
@@
  • $d_{TA}$: Impact of development tools on application security and functionality.
  • $d_{DD}$: Interconnections and potential bias propagation between datasets.
  • $d_{DM}$: Influence of data quality on model performance and fairness.
@@
  • $d_{MM}$: Dependencies between models, such as in transfer learning or model composition.
  • $d_{MA}$: Inheritance of model characteristics and potential vulnerabilities by applications.
@@
These dependencies not only form the core structure of the LLM supply chain but also serve as potential sources of security issues. Consequently, the scope of LLM supply chain security has expanded from individual models to encompass every component and dependency within the entire chain. This expansion significantly increases potential security risks, as the attack surface is no longer limited to a single model but covers the entire supply chain, including computational infrastructure, development tools, training data, and upstream models. More importantly, due to the interdependencies within the supply chain, security issues in any component or dependency can trigger a chain reaction, affecting all downstream elements and thus amplifying the impact of security incidents. For example, a vulnerability in computational resources ($d_{RM}$) could compromise the integrity of model training, leading to widespread model vulnerabilities that propagate to numerous applications ($d_{MA}$). Similarly, biases in a fundamental dataset ($d_{DM}$) could permeate through various models and ultimately affect the fairness of a wide range of applications. This complex dependency network underscores the importance of comprehensively considering supply chain security and provides a more systematic and holistic perspective for researching and addressing related security issues. Future research should focus on developing methodologies for tracking multi-level dependency relationships, assessing cascading impacts of vulnerabilities across different components, and implementing security measures that consider the entire LLM supply chain ecosystem, from computational infrastructure to end-user applications.
@@
Resources Source code, documentation, repository commits & license, package metadata, CI/CD configurations, etc. Large-scale datasets, pre-trained models, documentation, repository license, model cards, prompts, etc.
Dev Tools Version control systems (git etc.), IDEs, compilers, debuggers, third-party libraries, build systems, static analysis tools, etc. AI foundational frameworks (PyTorch, TensorFlow, etc.), third-party components & tools (RAG, vector database, quantization, etc.), model hubs (Huggingface etc.) etc.
Core Product Software applications, libraries, APIs, etc. Pre-trained models, datasets, model APIs chatbots, agents, etc.
Distribution Code repositories (GitHub, GitLab), package managers (npm, pip) Model hubs (HuggingFace, ModelScope), code repositories, AIOps cloud platforms
Deployment Software installation, cloud deployment, containerization (Docker) Model serving, API integration, containerized model deployment
Maintenance Feature updates, bug fixes, vulnerability patches, documentation updates Model retraining, fine-tuning, model drift monitoring, model patch, etc.
Quality Assurance Unit testing, code reviews, specification checks, fuzz testing, SAST, IAST, etc. Performance evaluation metrics, bias & privacy assessment, AI red teaming, etc.
@@
New in LLM Model clone detection, model dependency management, risk propagation (hallucination, bias & toxicity), model license compliance, model patch migration, prompt engineering best practices, model interpretability, etc.
@@
New in LLM Dataset poisoning, model deserialization attacks (e.g., backdoors), prompt injection, jailbreak attacks, vulnerable autonomous agents, code hallucination (squatting), model inversion attacks, membership inference attacks, adversarial examples, etc.

### v3

@@ Abstract.
The rapid advancement of large language models (LLMs) has revolutionized artificial intelligence, introducing unprecedented capabilities in natural language processing and multimodal content generation. However, the increasing complexity and scale of these models have given rise to a multifaceted supply chain that presents unique challenges across infrastructure, foundation models, and downstream applications. This paper provides the first comprehensive research agenda of the LLM supply chain, offering a structured approach to identify critical challenges and opportunities through the dual lenses of software engineering (SE) and security & privacy (S&P). We begin by establishing a clear definition of the LLM supply chain, encompassing its components and dependencies. We then analyze each layer of the supply chain, presenting a vision for robust and secure LLM development, reviewing the current state of practices and technologies, and identifying key challenges and research opportunities. This work aims to bridge the existing research gap in systematically understanding the multifaceted issues within the LLM supply chain, offering valuable insights to guide future efforts in this rapidly evolving domain.
LLM Supply Chain, Large Language Models
journal: TOSEMccs: Software and its engineering Open source modelccs: Software and its engineering Software libraries and repositoriesccs: General and reference Surveys and overviewsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
The landscape of artificial intelligence (AI) has been dramatically transformed by the rapid advancements in pre-trained large language models (LLMs) and multimodal large language models (MLLMs)  [^1] For simplicity in this text, both pre-trained LLMs and MLLMs will be collectively referred to as LLMs, and their supply chains will be referred to as the LLM supply chain in the subsequent sections.. Models such as GPT-4 (OpenAI, 2023), Gemini (Anil et al., 2023), and LLaMA (Touvron et al., 2023) have pushed the boundaries of what machines can understand and generate, exhibiting remarkable capabilities in processing and producing human-like text and multimodal content. These sophisticated models are trained on vast and diverse datasets, enabling them to perform a wide array of tasks ranging from natural language understanding and content generation.
As the adoption of LLMs continues to expand, the necessity for a robust and efficient supply chain to support their development, deployment, and maintenance becomes increasingly apparent. The LLM supply chain encompasses the lifecycle of these models, involving multiple stages such as data curation (e.g. clearnlab (Cleanlab, 2024) for data quality assurance and Hugging Face datasets (HuggingFace, 2024) for data management), model training (e.g. PyTorch Distributed (PyTorch, 2024) for distributed training), optimization (e.g. OmniQuant (OpenGVLab, 2024) for model quantization and mergekit (AI, 2024) for model merging), and deployment  (e.g. AutoGPT (Significant-Gravitas, 2024) for agent workflow orchestration and RAGFlow (InfiniFlow, 2024) for retrieval-augmented generation). This complex ecosystem requires seamless collaboration among various stakeholders, including model developers who design and refine architectures, data providers who supply necessary training data, infrastructure engineers who manage computational resources, and end-users who interact with final applications. Each stakeholder plays a critical role in ensuring the functionality and reliability of LLMs, and the interdependencies among them highlight the need for coordinated efforts to address challenges inherent in the supply chain.
@@
Current State & Challenges. Despite the significant progress and transformative potential of LLMs, the rapidly evolving landscape of their development and deployment presents many challenges across the entire supply chain. These challenges span multiple fileds, including security, privacy, software engineering, systems, networking, data science, and user experience design. In this paper, we focus specifically on Software Engineering (SE) and Security & Privacy (S&P). From the SE perspective, the unique characteristics of LLMs introduce a host of challenges that extend beyond traditional software development practices. The sheer scale and complexity of these models necessitate novel approaches to system design, development, and maintenance. Key challenges include model versioning and reproducibility (Ajibode et al., 2024), dependency management (Jiang et al., 2023), adapting continuous integration and deployment (CI/CD) practices for LLMs (Chen, 2024), comprehensive testing and quality assurance (Chang et al., 2024; Sun et al., 2024c), model interpretability and explainability (Singh et al., 2024), technical debt management (Bill Doerrfeld, 2024), and navigating the complex landscape of open-source licenses (Charlie Hull, 2023; Sriram Parthasarathy, 2024). From the S&P perspective, vulnerabilities have been identified at various layers of the LLM supply chain. These range from flaws in LLM development toolchains at the infrastructure layer (iumiro, 2023; Tom Bonner, 2023; Patrick Peng, 2024; Smartkeyss, 2024) to vulnerabilities (Eoin Wickens, Kasimir Schulz, 2024) or code poisoning (Zhao et al., 2024b) attacks in model hubs, and extend to remote code execution risks in downstream applications (Liu et al., 2023b). The pervasive nature of these vulnerabilities highlights the need for comprehensive security measures throughout the LLM supply chain. The potential for cascading effects due to model reuse and the difficulty in isolating and mitigating issues in complex LLM-based systems further exacerbate these challenges, often outstripping the establishment of robust security measures and software engineering practices.
@@
Research Gap & Our Work. Addressing these challenges necessitates a multidisciplinary approach that draws insights from fields such as software engineering, system architecture, security, and data governance. However, there exists a significant research gap in systematically tackling these multifaceted issues within the context of the LLM supply chain. To bridge this gap, we present the first comprehensive definition of the LLM supply chain, proposing a structured approach to analyze its challenges and opportunities across three key layers: infrastructure, foundation models, and downstream applications. We begin by establishing a clear definition of the LLM supply chain, encompassing its components and dependencies. Building upon this foundation, we examine each layer through the dual lenses of SE and S&P. For each layer, we present a vision for the robust and secure LLM supply chain, review the current state of practices and technologies, and identify key challenges and research opportunities. In the infrastructure layer, we explore issues related to computational resources, development tools, and datasets. For foundation models, we delve into the complexities of model training, testing, sharing, and maintenance. The downstream applications layer focuses on the integration and deployment of LLM-based systems in real-world scenarios. Through this comprehensive examination, we propose a research agenda that not only identifies current challenges but also anticipates future developments in the field. Our goal is to offer insights that can guide future efforts in this rapidly evolving domain, fostering innovation while upholding ethical standards and ensuring the responsible advancement of LLM technologies.
@@
  • Definition of the LLM supply chain: We are the first to present the definition of the LLM supply chain, which provides a foundational understanding of the LLM supply chain’s components and stages. The supply chain is delineated into three core layers: the model infrastructure layer (including computational resources, datasets, and toolchains for training, optimization, and deployment), the foundation model layer (encompassing training, testing, releasing, and ongoing maintenance), and the downstream application ecosystem (facilitating the integration of pre-trained models into various intelligent applications).
@@
  • Proposal of the Research Agenda: Based on the systematic review, we propose a forward-looking research agenda that addresses current challenges and anticipates future developments in the LLM supply chain. This agenda is designed to guide future efforts in the rapidly evolving domain of the LLM supply chain by suggesting multidisciplinary approaches that draw from SE and S&P. The goal is to foster advancements that maximize the transformative potential of LLMs while adhering to ethical standards and promoting sustainability.
@@ 2. Definitions and Significance
In this section, we start by defining the LLM supply chain (§ 2.1). We then explore three key aspects that highlight its significance: the novel challenges introduced by LLMs beyond traditional open-source software (OSS) (§ 2.2), the increased complexity of managing LLMs compared to conventional machine learning (ML) and deep learning (DL) models (§ 2.3), and the cascading implications beyond individual models or applications (§ 2.4). These elements collectively underscore the urgent need for a comprehensive research agenda focused on the LLM supply chain.
Figure 1. Definition and Each Component of the LLM Supply Chain.
@@ 2.1. Definitions
In the traditional domain of the OSS supply chain, numerous efforts (Ji et al., 2023c; Gao et al., 2024; He et al., 2020) have been made to define the scope and components of OSS supply chains. He et al. (He et al., 2020) defined the OSS supply chain as the end-to-end process of developing, releasing, deploying, and maintaining software from source code to final delivery to users. Ji et al. (Ji et al., 2023c) adopted a broader perspective, incorporating third-party components and defining the OSS supply chain as the network of supply relationships formed by upstream communities, source/binary packages, distribution markets, developers, maintainers, and foundations during software development and operation. Gao et al. (Gao et al., 2024) further delineated four types of open-source artifact reuse relationships (dependency installation, API calls, file copying, and code cloning), characterizing the OSS supply chain as the supply network arising from various code reuse practices across the software lifecycles.
Building upon these prior conceptualizations of OSS supply chains, we define the LLM supply chain as a complex network system $S=(C,D)$, where $C$ represents the set of components and $D$ represents the set of dependencies. This system involves all upstream data providers, model development communities, model repositories, distribution platforms, and app markets, as well as data scientists, researchers, engineers, maintainers, and end-users in the process of developing, distributing, and deploying pre-trained models. As illustrated in Figure 1, the LLM supply chain comprises five key components $C=\{R,T,D,M,A\}$: computational resources ($R$), tools ($T$), datasets ($D$), models ($M$), and applications ($A$). The first three components ($R$, $T$, and $D$) collectively form the infrastructure of the LLM supply chain.
@@
  • LLM Toolchain ($T$): Encompasses LLM development toolchains, including AI frameworks (e.g. PyTorch (PyTorch, 2024) and TensorFlow (TensorFlow, 2024)), dataset curation tools, model optimization tools (e.g. OmniQuant (OpenGVLab, 2024) for model quantization and mergekit (AI, 2024) for model merging), and downstream application frameworks (e.g. LangChain (LangChain-AI, 2024)).
@@
  • Datasets ($D$): Comprises large-scale text corpora (including natural language and code), specialized domain datasets, and multi-modal data collections.
@@
  • Models ($M$): Represents the various stages of LLM development, including architecture design, pre-training, fine-tuning, testing, releasing, deployment, and maintenance.
@@
  • Applications ($A$): Includes downstream LLM-based applications such as chatbots, autonomous agents, and domain-specific LLM solutions.
Within the LLM supply chain, there exist multiple complex dependencies arising from reuse and sharing of components across different stages. These dependencies can be denoted as $D=\{d_{ij}|i,j\in C\}$, where $C$ represents the set of components in the supply chain. Specifically, these dependencies are primarily driven by reuse and can be summarized as follows:
  • $d_{RM}$: Vulnerabilities in computational resources may lead to single points of failure, impacting model availability and security.
@@
  • $d_{TT}$: Interoperability and version compatibility issues between different development tools may introduce vulnerabilities or security flaws in the toolchain.
  • $d_{TD}$: Deficiencies in data processing tools can compromise dataset quality and integrity, affecting the security and fairness of models trained on these datasets.
  • $d_{TM}$: Reused tools and components in the development process may introduce vulnerabilities or unintended optimizations into models and applications.
@@
  • $d_{TA}$: The security and functionality of development tools can directly impact the security and reliability of LLM applications built using these tools.
  • $d_{DD}$: Interconnections and shared usage between datasets, either through direct reuse or derivation from common sources, can lead to the propagation of data quality issues.
  • $d_{DM}$: Data quality issues can severely impact the performance, security, and fairness of models trained on these datasets.
@@
  • $d_{MM}$: Dependencies between models, such as transfer learning or model merging, may facilitate the propagation of vulnerabilities or biases across different models.
  • $d_{MA}$: Applications may inherit characteristics, vulnerabilities, and biases from the models they employ, affecting their security and fairness performance.
@@
These dependencies not only form the core structure of the LLM supply chain but also serve as potential sources of security issues. Consequently, the scope of LLM supply chain security has expanded from individual models to encompass every component and dependency within the entire chain. This expansion significantly increases potential security risks, as the attack surface is no longer limited to a single model but covers the entire supply chain, including computational infrastructure, development tools, training data, and upstream models. More importantly, due to the interdependencies within the supply chain, security issues in any component or dependency can trigger a chain reaction, affecting all downstream elements and thus amplifying the impact of security incidents. For example, a vulnerability in computational resources could compromise the integrity of model training ($d_{RM}$), leading to widespread model vulnerabilities that propagate to numerous applications ($d_{MA}$). Similarly, biases in a fundamental dataset could permeate through various models ($d_{DM}$) and ultimately affect the fairness of a wide range of applications ($d_{MA}$). This complex dependency network underscores the importance of comprehensively considering supply chain security and provides a more systematic and holistic perspective for researching and addressing related security issues. Future research should focus on developing methodologies for tracking multi-level dependency relationships, assessing cascading impacts of vulnerabilities across different components, and implementing security measures that consider the entire LLM supply chain ecosystem, from computational infrastructure to end-user applications.
@@
Resources Source code, documentation, repository commits & license, package metadata, CI/CD configurations, etc. Large-scale datasets, pre-trained models, prompts, documentation, repository license, model cards, etc.
Dev Tools Version control systems (git etc.), third-party packages/libraries, IDEs, compilers, debuggers, build systems, static analysis tools, etc. AI foundational frameworks (PyTorch, TensorFlow, etc.), third-party components & tools (RAG, vector database, quantization, etc.), model hubs (Hugging Face, etc.), etc.
Distribution Code repositories (GitHub, GitLab, etc.), package managers (NPM, PyPI, etc.), online platforms  (Stack Overflow etc.) Model hubs (Hugging Face, ModelScope, etc.), dataset communities (Kaggle etc.), prompt hubs, code repositories, etc.
Maintenance Feature updates, bug fixes, vulnerability patches, documentation updates, etc. Model retraining, fine-tuning, model drift monitoring, model merging, etc.
Quality Assurance Unit testing, code reviews, specification checks, fuzz testing, SAST, IAST, etc. Performance evaluation, bias & privacy assessment, AI red teaming, etc.
@@
New in LLMs Model clone detection, model dependency management, risk propagation (hallucination, bias & toxicity), model license compliance, prompt engineering best practices, etc.
@@
New in LLMs Dataset poisoning, model deserialization attacks (e.g., backdoors), code hallucination, prompt injection, jailbreak attacks, vulnerable autonomous agents, etc.